Fields Company has two manufacturing departments, forming and painting. The company uses the weighted-average method of process costing. At the beginning of the month, the forming department has 37,000 units in inventory, 80% complete as to materials and 20% complete as to conversion costs. The beginning inventory cost of $84,100 consisted of $59,200 of direct materials costs and $24,900 of conversion costs. During the month, the forming department started 540,000 units. At the end of the month, the forming department had 40,000 units in ending inventory, 80% complete as to materials and 30% complete as to conversion. Units completed in the forming department are transferred to the painting department. Cost information for the forming department follows. Beginning work in process inventory $ 84,100 Direct materials added during the month 1,949,370 Conversion added during the month 1,254,270 Assume that Fields uses the FIFO method of process costing.

1. Calculate the equivalent units of production for the forming department.

2. Calculate the costs per equivalent unit of production for the forming department.
Answer #1

Unit transferred out = 37000+540000-40000 = 537000

Unit started and completed = 537000-37000 = 500000

a) Equivalent unit

Whole unit % Material EUP-Material % Conversion EUP-Conversion
Beginning inventory 37000 20% 7400 80% 29600
Unit started and completed 500000 100% 500000 100% 500000
Ending inventory 40000 80% 32000 30% 12000
Equivalent unit♥ 577000 539400 541600

b) Cost per equivalent unit

Material Conversion
Cost added 1949370 1254270
Equivalent unit 539400 541600
Cost per equivalent unit 3.61 2.32